Find the balance in the accoubt: $2,400 principal earning 2%, compounded quarterly, after 7 years
Answer by stanbon(75887)

Find the balance in the account: $2,400 principal earning 2%, compounded quarterly, after 7 years
------
A(t) = P(1+(r/n))^(nt)
------
A(7) = 2400(1+(0.02/4))^(4*7)
------
A(7) = 2400*1.1499
----
A(7) = $2759.69
=================
Cheers,
Stan H.